OpenStudy (anonymous):

11. What is the numeric relation between pyrimidine and purine bases in the DNA molecule? Is that relation valid in RNA molecules?

OpenStudy (alphadxg):

The DNA molecule is made of two bound polynucleotide chains that form a helical structure (the double helix). The binding of the two chains is between their nitrogen-containing bases and it always obeys the following rules: adenine (A), a purine base, binds with thymine (T), a pyrimidine base, and guanine (G), a purine base, binds to cytosine (C), a pyrimidine base. Therefore in one molecule of DNA there will be the same number of adenine (A) and thymine (T) and same number of cytosine (C) and guanine (G). The quantities of purine and of pyrimidine bases then will also be the same in a 50% proportion for each type. The relation A = T and C = G, or A/T = C/G = 1, is called Chargaff’s relation and the pairing rules described above are known as Chargaff’s rules.